UT 03 Sep 2021 - Night #2 of Restart




Night Staff:

OSA: S. Allanson

ISp: C. Bolyard

ISA: B. Rothberg, R.T. Gatto (Tucson secret Outpost)

Technical Wizard: J. Hill (Tucson)

Mtn Manager: P. Hartley

Sunset: 18:47

Twlight Ends: 20:07

Twilght Starts: 04:31

Sunrise: 05:51

Starting Config: LUCI-Bino (AGWs only)

Night Log:

UT 00:58 - AGW-1 and AGW-2 back up and operations - required software troubleshooting and some hardware checks on air pressure system and rotator thanks to S. Allanson and B. Bingham

UT 01:08 - powering up LBCs for possible use later, other GUIs for AGW/M3 work up

UT 01:13 - taking test biases with LBCs

UT 02:07 - opening dome

UT 02:15 - moving to pointing star

AGW-1/AGW-2 Checkout:

Pointing/Collimation:

UT 02:20 - Pointing completed

UT 02:24 - moving to nearby Persson standard to collimate on-axis

Guiding and SH WFS check

UT 02:29 - SH spots look good - cutout on leftside looks a little off, moving to slightly fainter (R=15.5mag) star to check contrast between star and background. We will refine the on-axis pupil position for the cutoff after the Pupil Wobble adjustment on M3. Present SX on-axis pupil position is 378.8, 244.8 (from post-processing a few dozen wfs images in IDL) Present DX on-axis pupil position is 376.6, 266.1 .

-visual check of the SH Guiding Hotspots - OK

SH-LUCIs_Check.jpg

AGW1/AGW2 Align Probe to Rotator Center

UT 02:43 - next step rotator center

SX - J. Hill
  • stopping guiding vis GCSGUI
  • slewing rotator to PA=0
  • setting AGW probe to nominal center: setxy -u 1 -x 0 -y 612.5
  • on-axis preset so focus is at nominal value
  • gcsclient readGuideCam left -e 1000 - this will fail on a 1 sec exposure, will need to take a 2nd at 90 sec
  • UT 02:55 gcsclient readGuideCam left -e 90000
  • slew to hold 360
  • /lbt/data/share/tcs/lbto.20210903/left_guider000001.fits
  • center at 113, 82
    x2 226 164
    detxy 241, 203
    compare hotspot 249,196
    0.40", 0.38"
  • UT 03:02 gcsclient readGuideCam left -e 90000 (for confirmation without any changes)
  • /lbt/data/share/tcs/lbto.20210903/left_guider000002.fits
  • center at 116, 80 (circle has some ellipticity)
    x2 232, 160
    detxy 247, 199 (officially close enough without changes)

DX - B. Rothberg
  • stopping guiding vis GCSGUI
  • slewing rotator to PA=0
  • ptincrement 1 CA -5 (1 for DX)
  • setting AGW probe to nominal center: setxy -u 2 -x 0 -y 612.5
  • on-axis preset so focus is at nominal value
  • gcsclient readGuideCam right -e 1000 - this will fail on a 1 sec exposure, will need to take a 2nd at 90 sec
  • gcsclient readGuideCam right -e 90000
  • slew to hold 360
  • /lbt/data/share/tcs/lbto.20210903/right_guider000001.fits
  • X=128, Y=114
  • correct for binning: X=256, Y=228
  • gethead DETSEC --> [18:526,16:510] add +17 X and +15 Y
  • X=263, Y=243
  • lbto@obs2:21 % gcsclient getHotspot right
    • Hotspot is (x,y): 270.00, 238.00
  • X=+7 Y=-5 ---> 1 pixel 0.05" --> X = +0.35" Y -0.25"
  • 0.6mm/" ---> X=+0.21mm, The ROT axis has approximately 68 counts/mm. X==>14.28 counts
  • Y should be in microns = Y= -0.15mm => Y = -150 microns
  • X--> ROT; Y-->RAD
    • Decreasing ROT moves toward larger X pixel
    • Decreasing RAD moves toward larger Y pixel --> so here we want to INCREASE RAD
  • from telescope@obs1, i ssh oac@oac
  • backup oacontrol.conf
  • edit oacontrol.conf --> find AGW2
    • increase homeoffset +14.28 counts
    • homeoffset=-33803 (from 33817)
    • increase ooffset 150 microns
    • oofset=419332 (from 41933)
  • update parameters with rdwrconfig -u 2
  • stop and restart GCS right
  • resend the active preset to the right
  • Now we remeasure!
    • Hotspot is (x,y): 270.00, 238.00
    • /lbt/data/share/tcs/lbto.20210903/right_guider000002.fits
    • Not sure what happened - image is just the star - no trail
    • #0004 seems ok - but very tight circle
    • X=126, Y=112
    • Conversion: X=269, Y=234
  • DONE!

Pupil Wobble to align M3 for LFBG / RFBG

  • Data Collection on Wiki
  • UT 03:30 - Active preset to R=13.1 mag star
  • running in parallel IDL windows, data directories: /home/lbto/data/20210903_lfbg & /home/lbto/data/20210903_rfbg
  • collection is by lbto@obs2 (Barry), analysis is by telescope@robs (John)
  • UT 03:36 - running pupil_collect #1 on both sides (hit stop WFSing on both sides after I executed IDL script)
    • timed OK for left, had to restart right (it complained WFSing was going on!)
    • plot with /lbt/lbto/supportscripts/MetrologySupport/m3grace.py -s SX
  • UT 03:52 Apply +100 Tip on both sides
    /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s DX -p 100 -t 0
  • UT 03:54 - running pupil_collect #2 (both wobbles got smaller)
  • UT 04:08 Apply +50 Tip on both sides
  • UT 04:10 - running pupil_collect #3
  • UT 04:24 Apply -20 Tip -50 Tilt on SX , Apply +50 Tip on DX
  • UT 04:25 - running pupil_collect #4
  • UT 04:38 Apply 0 Tip -50 Tilt on SX , Apply -20 Tip +50 Tilt on DX
  • UT 04:43 - running pupil_collect #5
  • UT 04:57 Apply 0 Tip +50 Tilt on SX to undo, Apply 0 Tip +25 Tilt on DX
  • UT 04:59 - running pupil_collect #6
  • UT 05:10 Apply nothing to SX, Apply -35 Tilt on DX and adopt the totals

ABS SX -90 Tip -320 Tilt ABS DX -160 Tip +120 Tilt LFBG:
  • LFBG Pupil Wobble Analysis Plot 20210903:
    LFBG_20210903_wobble.jpg
/lbt/lbto/supportscripts/MetrologySupport/m3grace.py script version of 02-Sep-2021
ls -1t *.wobble

filename               Xavg     Yavg     Radius    Angle1 (rms)     Angle2 (rms)   pts
pupil_04:58:09.wobble  381.122  245.365    1.007   43.175 ( 93.2)  122.092 ( 43.4) 6
pupil_04:42:09.wobble  381.352  245.083    1.738    1.915 ( 95.6)  127.935 ( 18.5) 6
pupil_04:25:04.wobble  381.423  245.563    0.866   -0.867 ( 95.9)   48.339 (140.0) 6
pupil_04:09:51.wobble  381.382  245.177    0.827  -22.181 ( 83.1)  -51.161 (135.4) 6
pupil_03:53:00.wobble  381.237  245.242    0.871   -3.711 ( 80.6)   64.352 ( 21.6) 6
pupil_03:35:55.wobble  381.233  244.987    2.355   44.857 ( 93.8)   84.107 ( 15.7) 6

RFBG:
  • RFBG Pupil Wobble Analysis Plot 20210903:
    RFBG_20210903_wobble.jpg
/lbt/lbto/supportscripts/MetrologySupport/m3grace.py script version of 02-Sep-2021
ls -1t *.wobble

filename               Xavg     Yavg     Radius    Angle1 (rms)     Angle2 (rms)   pts
pupil_04:58:23.wobble  379.912  265.320    2.126   50.559 (106.2)   60.043 ( 24.9) 6
pupil_04:43:57.wobble  378.762  265.185    1.379   63.055 ( 90.2)   60.899 ( 17.2) 6
pupil_04:25:24.wobble  379.470  264.868    1.929  -53.772 ( 80.7)   46.021 ( 30.3) 6
pupil_04:09:54.wobble  379.378  264.582    2.029  -39.680 ( 92.2)   41.867 ( 15.4) 6
pupil_03:53:32.wobble  379.530  264.133    2.735   -3.433 (100.1)   55.867 ( 30.9) 6
pupil_03:37:00.wobble  379.395  265.035    3.152   42.958 ( 95.4)   73.748 ( 11.9) 6

JMH updated the M3 Tip-Tilt positions for M3 the following afternoon:
/lbt/data/config/tcs/OSS/positions.conf

the order is:  tip   tilt  piston rotation

grep 20210903 positions.conf -A 2 -B 3

# SX M3 LFBG positions 20201227 JMH/OPK
#oss.side[0].terc.focal_stations.bentGregorianFront             map  0:-220.0   1:-270.0  2:0.0  3:+28.000
#oss.side[0].terc.focal_stations.bentGregorianArgos             map  0:-220.0   1:-270.0  2:0.0  3:+28.007
# SX M3 LFBG positions 20210903 JMH/BR
oss.side[0].terc.focal_stations.bentGregorianFront              map  0:-90.0   1:-320.0  2:0.0  3:+28.000
oss.side[0].terc.focal_stations.bentGregorianArgos              map  0:-90.0   1:-320.0  2:0.0  3:+28.007
--
# DX M3 positions for 20201227 RFBG JMH/OPK
#oss.side[1].terc.focal_stations.bentGregorianFront             map  0:-340.00  1:80.00 2:0.0  3:-25.727
#oss.side[1].terc.focal_stations.bentGregorianArgos             map  0:-340.00  1:80.00 2:0.0  3:-25.734
# DX M3 positions for 20210903 RFBG JMH/BR
oss.side[1].terc.focal_stations.bentGregorianFront              map  0:-160.00  1:120.00 2:0.0  3:-25.727
oss.side[1].terc.focal_stations.bentGregorianArgos              map  0:-160.00  1:120.00 2:0.0  3:-25.734

ADD PUPIL WOBBLE PLOTS HERE


Field Aberration Data Collection for LFBG / RFBG

UT 05:15 - pointing check

UT 05:18 - Active Preset BS9188

UT 05:19 - Guide Preset on-axis to BS9188
  • Seeing 0.66" SX, 0.56" DX
  • field_collect,/IDL - turns off the WFS
  • starlist BS9188_field.list
  • didn't crash yet!
  • now it crashed:
  • IRCR: preset, BS9188, MODE=GUIDE, GS=0, PA=61.6590 (iif_preset)

    % Compiled module: IRC_PRESETTELESCOPE.
    => right preset complete
    % Compiled module: COLLIMATE_MULTIPLE.
    % Ambiguous keyword abbreviation: SEND.
    % Execution halted at: FIELD_COLLECT 214 /home/lbto/idl/wfsc/field/field_collect.pro
    % $MAIN$

  • UT 05:23 - Dr Doug is on the case....downtime
  • Dr. Doug fixed the IDL request - routne is working
  • DIMM @zenith=0.95"
  • #0040 - Right side SH spots only bottom 4 visible, didn't see this for #0040 on the left side
  • DX/AGW-2 on-axis full illumination of SH, off-axis SH spots are not illuminated
    • Rotator angle=209 - again showing shadowed SH spots (#0046) (was at PA=114.616), gs 12
  • UT 05:49 - exited out of right side
  • Left is continuing to collect data
  • Completed UT 06:28 - seeing was <1" the whole time
Need to troubleshoot DX AGW2 shadow off-axis
  • send Active mode preset to 9188, gs=1 first, then proceed to 2
  • IDL> iif_preset,BS9188,mode=active,gs=1,/yaxis
# Target star: 0 BS9188
23:30:33.1379 +38:18:57.763 RADEC 2000.0 J2000 -10.3 -2.4 12.42 R 0.87 B_V 1.646
# Guide star: 1
23:30:35.9665 +38:19:48.459 RADEC 2000.0 J2000 -0.8 -7.4 13.62 R 0.96 B_V 0.630

IRCR: preset, BS9188, MODE=ACTIVE, GS=1, PA=33.1857 (iif_preset)
  • no vignetting - x=-0.218, y=36.211

  • IDL> iif_preset,BS9188,mode=active,gs=2,/yaxis

# Target star: 0 BS9188
23:30:33.1379 +38:18:57.763 RADEC 2000.0 J2000 -10.3 -2.4 12.42 R 0.87 B_V 1.646
# Guide star: 2
23:30:41.1801 +38:17:54.419 RADEC 2000.0 J2000 13.6 11.3 9.99 R 1.99 B_V 0.630

IRCR: preset, BS9188, MODE=ACTIVE, GS=2, PA=123.793 (iif_preset)
  • no vignetting, x=0.275, y=68.198

  • IDL> iif_preset,BS9188,mode=active,gs=4,/yaxis

# Target star: 0 BS9188
23:30:33.1379 +38:18:57.763 RADEC 2000.0 J2000 -10.3 -2.4 12.42 R 0.87 B_V 1.646
# Guide star: 4
23:30:28.8256 +38:16:21.809 RADEC 2000.0 J2000 -10.8 -14.6 13.89 R 1.14 B_V 0.630

IRCR: preset, BS9188, MODE=ACTIVE, GS=4, PA=-161.995 (iif_preset)
Rotator =295 degrees
  • definite vignetting, x=-0.179, y=98.354
GS4_BS9188_Vignetting.jpg
  • IDL> iif_preset,BS9188,mode=active,gs=8,/yaxis

# Target star: 0 BS9188
23:30:33.1379 +38:18:57.763 RADEC 2000.0 J2000 -10.3 -2.4 12.42 R 0.87 B_V 1.646
# Guide star: 8
23:30:25.3919 +38:15:44.181 RADEC 2000.0 J2000 2.3 -11.1 15.00 R 1.09 B_V 0.630

IRCR: preset, BS9188, MODE=ACTIVE, GS=8, PA=-154.723 (iif_preset)
  • Definite 2x+ more vignetting, x=0.199, y=128.192
GS8_BS9188_Vignetting.jpg

  • Moving to opposite side of the field to check if this is rotator related
  • IDL> iif_preset,BS9188,mode=active,gs=11,/yaxis

# Target star: 0 BS9188
23:30:33.1379 +38:18:57.763 RADEC 2000.0 J2000 -10.3 -2.4 12.42 R 0.87 B_V 1.646
# Guide star: 11
23:30:50.2878 +38:20:46.628 RADEC 2000.0 J2000 0.6 -13.7 15.00 R 1.02 B_V 0.630

IRCR: preset, BS9188, MODE=ACTIVE, GS=11, PA=61.6590 (iif_preset)
Rotator = 202 degrees

  • Definite vignetting, appears the same as before - so unlikely a rotator contribution
GS11_BS9188_Vignetting.jpg

This issue is IT 8457 CRITICAL The problem was confirmed to be the AGW2 cover upside down the following day.

On-Axis Pupil Positions

JMH used the wfs images collected by field_collect to measure the on-axis pupil positions (after the M3 alignment. The processed files are in /home/lbto/data/202109_LUCI/ based on wfs images in /home/lbto/data/20210903/Xfbg_field/.

The SX on-axis pupil position is 380.9, 244.6 . WFScam_hotspot_x/y was updated in LUCI_L.cfg the following evening. The DX on-axis pupil position is 378.9, 264.9 (measured with fewer pupils, and indistinguishable from the previous position).


MODS AGW 5&6 Checkout:

UT 06:29 - reconfiguring for MODS

UT 06:23 - MODS-2 found in weird state with all services (except lbttcs) disconnected, brought up as many servies as possible. Could not restart M2RC. Only need AGW6. Envt services were restarted. IT 8458 HIGH

UT 06:50 - pointing check AGW 5/6

UT 06:59 - had weird error where some IDL process was hijacking the authorization

UT 07:00 - moving to R=15.9 star to visually check SH spots

AGW6 seems to pick different stars than AGW5 in crowded fields

SH spots look OK

MODS AGW5/6 Check on Probe w.r.t Rotator Center

UT 07:06 - guideprobe check w/rotator center
SX - J. Hill
  • Stop guiding
  • Slew to hold 0 LDG
  • ptincrement 0 CA -5
  • setting gprobe 0 0 in MODS GUI command (BR did this for both MODS)
  • gcsclient readGuideCam left -e 1000 (throwaway exposure)
  • gcsclient readGuideCam left -e 90000
  • slew to hold 360 LDG
  • lbt/data/share/tcs/lbto.20210903/left_guider00000X.fits
  • center 183,171
    detxy 258, 192
    hotspot 254, 201 (declared to be close enough ~1 arcsec)

DX - B. Rothberg
  • Stop guiding
  • Slew to hold 0 RDG
  • ptincrement 1 CA -5
  • setting gprobe 0 0 in MODS GUI command (doing this for both MODS)
  • gcsclient readGuideCam right -e 1000 (throwaway exposure)
  • gcsclient readGuideCam right -e 90000
  • slew to hold 360 rdg
  • lbt/data/share/tcs/lbto.20210903/right_guider000005.fits
  • X=183, Y=165
  • Conversion:
    • DETSEC [85:449,51:424]
    • X=183+84 = 267, Y = 165+50 = 215
    • Hotspot is (x,y): 267.00, 240.00
    • Y differential is 35 pixels - this is what we expected (can be as much as 40 pixels in Y)

Field Aberration Collection for MODS at LDG / RDG

UT 07:23 - moving to Stone P1 (since BS9188 will be transiting at > 80 deg elevation shortly)

UT 07:24 - Acquire to check pointing

UT 07:26 - Utility Chiller high pressure alarm went off S. Allanson waking P. Hartley

UT 07:31 - Active preset for collimation

UT 07:32 - Guide preset to do the field collection

These directories are:
  • data/20210903/ldg_field
  • data/20210903/rdg_field

UT 07:35 - starting data collection (note: Stone P list is Stone_P1.list everything else is *_field.list)

UT 07:56 Completed 4 stars worth of field aberration data for MODS - closing the hatch of each MODS.
Analysis of this data is reported in Issue 8315, and it does show general agreement with the larger data sets taken at LFBG / RFBG.

UT 07:57 - reconfigure for LBCs

LBC Checkouts


Superfoc Sequence

Field 23h00m+35d

0,180, then 270, 90 OBs

UT 08:15 - sending preset to CP field for collimation

UT 08:18 - dohybrid



COPOINTING: B=82837 R=82833
Pointing updates: delta_IE = 6.92", delta_CA = 6.63"
Mirror updates: dX(mm) dY(mm) dRX(") dRY(")
SX: 0.56 0.24 5.14 -12.07
DX: -0.63 -0.28 -6.01 13.51
-------------------------------------------------------------------------

Elevation 71-72 degrees
UT 08:35 - running PA 0 superfoc

UT 08:45 - running PA 180 superfoc

UT 08:51 - running PA 90 superfoc

UT 08:59 - running PA 270 superfoc
Rotator Center

Using 20h60d field first

UT 09:07 - sending CP for collimation

Airmass 1.51 (el 41)

pa0.ob has a correct Dec, the other three have Dec=-90, just rotates the instrument

UT 09:09 - running dofpia

UT 09:23 - sending copoint script

lbcrangebal Is NOT working on OBS2 it cannot read a critical dat



COPOINTING: B=92412 R=92408
Pointing updates: delta_IE = -4.27", delta_CA = 3.67"
Mirror updates: dX(mm) dY(mm) dRX(") dRY(")
SX: 0.61 0.21 4.43 -13.18
DX: -0.62 -0.18 -3.95 13.22
-------------------------------------------------------------------------
UT 09:38 - running pa0.ob

15min downtime for lbcrangebal not functioning on OBS2 - this is unacceptable as ROBS is not a stable setup

complains it cannot find "Rotcen.dat"

UT 09:41 - running pa180.ob

Airmass 1.645

UT 09:43 - running pa90.ob

UT 09:45 - running pa270.ob

Using 01h50d field

UT 09:47 - sending CP script for collimation

UT 09:49 - running dofpia

UT 09:54 - CP script



COPOINTING: B=95448 R=95445
Pointing updates: delta_IE = 0.25", delta_CA = 2.26"
Mirror updates: dX(mm) dY(mm) dRX(") dRY(")
SX: 0.62 0.32 6.87 -13.33
DX: -0.62 -0.27 -5.80 13.29
-------------------------------------------------------------------------
UT 09:57 - sending pa0.ob

Airmass 1.048 (el 72)

UT 09:58 - sending pa180.ob

LBC-Blue 095937 - missing segments of the CCD

too few downloaded pixels - this corresponds to PA=0

UT 10:02 - -rerun PA0

UT 10:04 - sending pa90.ob

UT 10:07 - sending pa270.ob
Science-like OB : Tadpole HCG98

UT 10:10 sending CP script for collination

UT 10:24 - bouncing around in z4 for red - hit shift+S

UT 10:25 - sending CP for HCG98



COPOINTING: B=102553 R=102549
Pointing updates: delta_IE = -16.63", delta_CA = 7.11"
Mirror updates: dX(mm) dY(mm) dRX(") dRY(")
SX: 0.60 0.36 7.72 -12.82
DX: -0.60 -0.33 -6.98 12.93
-------------------------------------------------------------------------
UT 10:27 - sending HCG98 OB (lets see if we get further than 1 dither!)

Airmass 1.377

UT 11:12 - completed 28/30 bino exposures

Non-Sidereal Observation : Comet 67P/Churyumov-Gerasimenko

(find in NSIGUI by searching for "Churyumov-Gerasimenko", and then selecting one of the numerical choices out of the log. 90000693_2021-Sep-03.txt )

UT 11:13 - co pointing for Comet 67P/Churyumov-Gerasimenko

UT 11:15 - run dofpia

- shift+S - doing the same z4 oscillations for both Red and Blue

UT 11:24 - co-pointing script

lbcrabgebal error

Observe> lbcrangebal
Timestamp of LBC-Blue image to process. (112532):
Timestamp of LBC-Red image to process. (112529):
Position cursor on Pointing Star and hit the Space Bar...
-2
ERROR: Attempt to access undefined local variable `XsrcB'.

"lp ("On LBCB image, star is at pixel coordinates: "//XsrcB//", "//Ys ..."
line 303: Observe$lbcrangebal.cl
called as: `lbcrangebal ()'
Observe>
restarted IRAF same error

now closing out RB_Science and reloading



COPOINTING: B=112532 R=112529
Pointing updates: delta_IE = 0.81", delta_CA = -12.49"
Mirror updates: dX(mm) dY(mm) dRX(") dRY(")
SX: 0.63 0.25 5.37 -13.59
DX: -0.56 0.04 0.86 12.09
-------------------------------------------------------------------------

Send offsets to the telescope?
Answer 'y' or 'Y' to send, anything else aborts this script.
sendT is now 0

Done!
didn't give me option for y or n

Now it works....



COPOINTING: B=112532 R=112529
Pointing updates: delta_IE = 0.32", delta_CA = -11.90"
Mirror updates: dX(mm) dY(mm) dRX(") dRY(")
SX: 0.62 0.23 5.03 -13.26
DX: -0.55 0.06 1.20 11.77
-------------------------------------------------------------------------

NSI GUI override - sending script

60 seconds 4 filters (BVRI)

Comet 67P:Churyumov-Gerasimenko_V3.jpg

UT 11:53 - finished for the night

Closing up

UT 11:55 - powering off LBCs

TMS was run in passive mode during LBC checkout from about 0800 UT on 20210903. OPK collected the logs from TMS and from FPIA in telescope@robs:/home/telescope/TMS_logs/20210903.

http://people.lbto.org/~cveillet/IQ/20210903/


Proposed Plan of Action:

  • LUCIs
  • Telescope Infrastructure
    • DONE SX M3 Alignment - Critical (3 hrs) - requires a LUCI AGw - both sides can be done in parallel
      • Pupil alignment check - find a star and center it - need Bent Gregorian (1 hr) - need to do before AGW checkout, but active optics needs to be working
      • Pupil wobble (2 hrs) - requires LUCI-1/AGW-1 to guide, collimate, SH spots well illuminated, and hotspot centered - can be done before AGW Rot ZP and Transform
    • DONE DX M3 Alignment - Critical (3 hrs) - requires a LUCI AGw- both sides can be done in parallel
      • Pupil alignment check - find a star and center it - need Bent Gregorian (1 hr) - need to do before AGW checkout,but active optics needs to be working
      • Pupil wobble (2 hrs) - requires LUCI-2/AGW-2 to guide, collimate, SH spots well illuminated, and hotspot centered - can be done before AGW Rot ZP and Transform
  • MODS
    • DONE Basic Checkout
      • Verify collimation
      • Verify AGW 5/6 Hotspots are centered
    • MODS-2 only:
      • MODS Rotator Center measure (use M71 at beg of night or M52 later) - requires pointing mask (40" square field stop & dot-circle with radius 45") - high (0.5hr)
      • Spectrophotometric standard
        • all modes medium (if clear)
          • Grating
          • Prism
        • dual grating high (if some clouds)
          • shake out of system
      • Test Science OB
        • Dice5 imaging
        • Sidereal spectroscopy
        • Non-Sidereal spectroscopy (low)
  • Contingency #1 (if LUCI AGWs and MODS AGWs not available)
    • LBC
      • DONE Collimation - DoHybrid, then DOFPIA (~1hr)
      • DONE Superfocus Sequence, PA=0, 180 (1hr)
      • DONE Rotator Center Checks (pairs of 0,180; 90, 270) ~0.5hr
      • DONE Test Science OB ~ 1hr
        • Sidereal
        • Non-Sidereal
      • Photometric Fields and Spectrophotometric Stars (to establish ZPs) ~ 1-1.5hrs
        • Requires multiple airmass visits (clear conditions, any seeing)
        • Cycle through all filters
        • This is now a high priority because we need to define ZPs
      • Twilight Flats
  • Contingency #2
    • PEPSI
      • PFU alignment
      • PEPSI pointing model (poor weather)
    • LUCI-1 restart activities (if MODS, LBCs, and PEPSI PFU can't be used)
      • Check Rotator Center
      • rotator zeropoint (requires MOS mask) - (0.5 hr)
        • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
      • Imaging check
        • dither in field (0.5hr) -
        • dither out of field (0.5hr) - M52 K and some J-band

Current Status

  • LBCs
    • LBC-Blue - available
    • LBC-Red - available
  • MODS offline
    • MODS-1 - offline - (disk issues with M1RC - unable to find boot disk) - cooling has not started
      • AGW-5 - available
    • MODS-2 - OK (IMCS lock check TBD)
      • AGW-6 - available
  • LUCI-1 - available
    • AGW-1 available
  • LUCI-2 offline
    • ´╗┐AGW-2 available
  • PEPSI
    • AGW-7/8 available
    • Software environmental system issues

Closed Dome Activities

  • Test Preset before opening (MODS)
  • ARGOS Cal-Unit test for M3 alignment and pupil wobble checkout - postponed
  • LUCI-1 Only Calibrations
    • Flexure data for G150
      • /home/dthomspons/Luci/FlexureCompensation
      • README - Calib15 - 2hrs
      • Ks filter, Argon lamp, optical sieve mask
      • take dark frame
      • Manual Control MCSGUI for elevations
    • Wavelength data for G150
      • G150 Ks, Argon, take image through slit, then put in grating, line should show up where the slit is located (within a few pixels of the image of the slit)
      • Check a few lines (2 or 3)
  • MODS Calibrations
    • Imaging Flats
      • Dual
      • Red
      • Blue
    • Spectroscopy (flats, slitless flats, arcs)
      • Dual
      • Red
      • Blue
    • Bias frames
      • 2x1 and 2x2 bias
      • 1x1 Binned Biases
    • Focus Scripts:
-- BarryRothberg - 02 Sep 2021

  • TMS plot from SX side:
    tms_left_20210903_telescope.jpg

  • TMS plot from DX side:
    tms_right_20210903_telescope.jpg
I Attachment Action Size Date Who Comment
Comet 67P:Churyumov-Gerasimenko.jpgjpg Comet 67P:Churyumov-Gerasimenko.jpg manage 116 K 03 Sep 2021 - 11:40 BarryRothberg  
Comet 67P:Churyumov-Gerasimenko_V2.jpgjpg Comet 67P:Churyumov-Gerasimenko_V2.jpg manage 116 K 03 Sep 2021 - 11:42 BarryRothberg  
Comet 67P:Churyumov-Gerasimenko_V3.jpgjpg Comet 67P:Churyumov-Gerasimenko_V3.jpg manage 108 K 03 Sep 2021 - 11:45 BarryRothberg  
GS11_BS9188_Vignetting.jpgjpg GS11_BS9188_Vignetting.jpg manage 130 K 03 Sep 2021 - 06:11 BarryRothberg  
GS4_BS9188_Vignetting.jpgjpg GS4_BS9188_Vignetting.jpg manage 101 K 03 Sep 2021 - 06:05 BarryRothberg  
GS8_BS9188_Vignetting.jpgjpg GS8_BS9188_Vignetting.jpg manage 110 K 03 Sep 2021 - 06:05 BarryRothberg  
LFBG_20210903_wobble.jpgjpg LFBG_20210903_wobble.jpg manage 49 K 04 Sep 2021 - 08:06 JohnHill LFBG Pupil Wobble Analysis Plot 20210903
RFBG_20210903_wobble.jpgjpg RFBG_20210903_wobble.jpg manage 38 K 04 Sep 2021 - 08:07 JohnHill RFBG Pupil Wobble Analysis Plot 20210903
SH-LUCIs_Check.jpgjpg SH-LUCIs_Check.jpg manage 225 K 03 Sep 2021 - 05:17 BarryRothberg  
tms_left_20210903_telescope.jpgjpg tms_left_20210903_telescope.jpg manage 351 K 06 Sep 2021 - 09:29 JohnHill TMS plot from SX side
tms_right_20210903_telescope.jpgjpg tms_right_20210903_telescope.jpg manage 325 K 06 Sep 2021 - 09:30 JohnHill TMS plot from DX side
Topic revision: r13 - 10 Sep 2021, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback